# CAGE: Curvature-Aware Gradient Estimation For Accurate Quantization-Aware Training

### v1

@@ 3.2 Theoretical Analysis
$x_{t+1}=x_{t}-\alpha(\widetilde{\nabla}f(x_{t})+\lambda(x_{t}-Q(x_{t}))),\quad t=0,1,\dots.$ (3)
We use the following assumptions regarding the smoothness of loss and nature of stochastic noise.
@@
Both assumptions are quite standard in the optimization literature and, in some sense, minimal to get meaningful convergence guarantees. Additionally, to handle quantization operator $Q$ in the analysis, it is common to approximate the actual non-smooth quantization operator $Q$ with smooth surrogates with annealing hyperparameters.
@@
$\mathbb{E}\|\nabla_{\rm\lambda P}f(Q(\hat{x}))\|^{2}\leq\frac{2}{\sqrt{T}}\left(f(x_{0})-f^{*}+\lambda\max_{x\in[x_{0},x_{T}]}\|x-Q(x)\|\|x_{0}-x_{T}\|+\frac{L\sigma^{2}}{2}\right)\max(1,\tfrac{L}{\sqrt{T}}),$
where $\hat{x}$ is a random iterate drawn from $\{x_{0},x_{1},\dots,x_{T-1}\}$ uniformly at random.
@@ Discussion.
First, observe that the convergence bound shows strong ergodic convergence to a Pareto-optimal state of problem 1, with a convergence rate that is optimal in the unquantized case. The convergence measure is defined with respect to our proposed Pareto-gradients (2) and, importantly, there is no non-vanishing term in the upper bound because of quantization error. In practical settings, both the distance $\|x_{0}-x_{T}\|$ and the quantization error $\|x-Q(x)\|$ are expected to remain bounded for any total number of training steps $T$. Therefore, this implies that the iterates (3) achieve $\mathcal{O}(\frac{1}{\sqrt{T}})$ ergodic convergence to the Pareto-optimal solution.
@@ 3.3 Practical Implementation
Since our main application is in training and fine-tuning LLMs, we implement CAGE on top of AdamW [24]. In contrast to LOTION [18], we focus on the decoupled update rule and keep the correction term outside the preconditioning path (Alg. 1), which we have found to work better in practice for this application. The base optimizer AdamW processes the mini-batch gradients $g_{t}$ obtained from STE, and the CAGE correction acts as a lightweight, elementwise post-step on the parameters. This decoupling has multiple practical advantages: (1) it preserves the optimizer’s well-understood behavior, (2) it keeps the Pareto stationarity direction xQ(x)x-Q(x) from being distorted by preconditioning, which empirically stabilizes training in low-bit regimes, and (3) it avoids numerical errors by decoupling terms with different behavior.
@@
4:Initialize $m_{0}\leftarrow 0,\;v_{0}\leftarrow 0,\;\widehat{e}_{0}\leftarrow 0$
@@
10:  $\lambda_{t}\leftarrow\lambda\cdot\dfrac{r_{t}-s}{1-s}$
11:  Sample minibatch and compute stochastic gradient $g_{t}$ with quantized forward pass.
12:  $x_{t}\leftarrow(1-\alpha\omega)\,x_{t}$ $\triangleright$ decoupled weight decay
13:  $m_{t}\leftarrow\beta_{1}m_{t-1}+(1-\beta_{1})g_{t}$
14:  $v_{t}\leftarrow\beta_{2}v_{t-1}+(1-\beta_{2})\,g_{t}\odot g_{t}$
15:  $\hat{m}_{t}\leftarrow m_{t}/(1-\beta_{1}^{t})$; $\hat{v}_{t}\leftarrow v_{t}/(1-\beta_{2}^{t})$
16:  $\tilde{x}_{t+1}\leftarrow x_{t}-\alpha\,\hat{m}_{t}/\left(\sqrt{\hat{v}_{t}}+\varepsilon\right)$
17:  $e_{t}\leftarrow x_{t}-Q(x_{t})$ $\triangleright$ quantization error (no grad)
18:  $x_{t+1}\leftarrow\tilde{x}_{t+1}-\alpha\,\lambda_{t}\,e_{t}$ $\triangleright$ decoupled correction
19:end for
20:return $x_{T}$
@@
$\tilde{x}_{t+1}=x_{t}-\alpha\cdot\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\varepsilon}\quad\text{with}\quad\hat{m}_{t}=\frac{m_{t}}{1-\beta_{1}^{t}},\;\hat{v}_{t}=\frac{v_{t}}{1-\beta_{2}^{t}}$
@@
$e_{t}\coloneq x_{t}-Q(x_{t}),\qquad x_{t+1}=\tilde{x}_{t+1}-\alpha\lambda_{t}\bar{e}_{t}$
@@
$r_{t}\coloneq\tfrac{t}{T},\qquad\lambda_{t}=\begin{cases}0,&r_{t}\leq s,\\ \lambda\cdot\dfrac{r_{t}-s}{1-s},&r_{t}>s.\end{cases}$
In practice, we have found parameters $s\in[0.8,0.9]$ with a linear ramp and $\lambda\in[0.5,2]$ work robustly; we will ablate these choices in §4. The ramp prevents abrupt shifts in the optimization objective and reduces loss spikes in training.
@@ 4 Experiments
Figure 1: Validation loss versus model size (bytes) for weight-only quantization using W2/W3/W4, comparing CAGE+QuEST to baseline QuEST and BF16. Observe that CAGE yields consistently lower loss.
Figure 2: Fitted parameter efficiency $\mathrm{eff}(P)$ across precisions: CAGE increases effective model capacity over QuEST, peaking near 4-bit.
@@
Our experiments use a row-wise quantizer as in QuEST [26]. Let $x\in\mathbb{R}^{d}$ denote a tensor row , and let $H\in\{\pm\tfrac{1}{\sqrt{d}}\}^{d\times d}$ be a (Walsh-)Hadamard transform. We define the pre-quantized tensor
@@
For bit-width $b$, we quantize symmetrically with integer bounds $q_{\max}=2^{b-1}-1$ and $q_{\min}=-2^{b-1}$. Using the pre-computed MSE-optimal Gaussian clipping factor $k_{b}$ (per bit-width), with $\sigma=\sqrt{\tfrac{1}{d}\sum_{i}z_{i}^{2}}$, we set the scale
$s=\frac{k_{b}\,\sigma}{q_{\max}},$
then apply
@@
Note that CAGE itself is quantizer-agnostic. It augments the optimizer update via the instantaneous quantization error $e_{t}=x_{t}-Q(x_{t})$ (see §3.3). We choose the QuEST quantizer here because it is a strong SOTA baseline for low-bit training. Other quantizers can be used without changing CAGE.
Pretraining experiments.
We test our method by pre-training Llama-style transformers with parameter counts $N\in\{30\text{M},50\text{M},100\text{M},200\text{M},800\text{M}\}$ under QAT with weight/activation bit-widths $b\in\{2,3,4\}$, using QuEST-style pre-quantization transforms as our baseline pipeline [26]. Training uses C4 with a fixed budget of $D=100\times N$ tokens (TPP), where $D$ is the number of tokens and $N$ is the number of parameters in the model. Master weights are stored FP32, aupdates follow AdamW with decoupled weight decay[24] as in 1. Experiments are done on 8xH100 GPUs and common hyperparameters (LR schedule, warmup, clipping, etc.) are in Appendix B. Each pretraining average is done over three seeds and the mean and standard deviation are reported in Table 1. Models up to $200M$ are visualized in Figure 2, although the standard deviation is too small to be visible in Figure 2.
We compare CAGE against (i) QuEST with Hadamard outlier mitigation (our strong baseline) and (ii) a high-precision BF16 reference. Figure 2 shows validation loss vs. model size for $b\in{2,3,4}$, with CAGE, QuEST baseline, and BF16 reference values.
@@
Method 30M 50M 100M 200M 800M
CAGE + HT 26.277 22.747 18.944 16.166 12.049
QuEST + HT 26.475 23.062 19.123 16.311 12.203
CAGE (no HT) 27.287 23.781 19.630 16.596 13.089
QuEST (no HT) 27.401 23.991 19.799 16.701 13.852
BF16 (reference) 24.715 21.491 17.923 15.422 11.541
Scaling law fitting.
To compare methods beyond pointwise losses, we fit a separable data/model/precision-scaling law to the validation loss. Following references [14; 26; 16], we fit a law of the form:
@@
where $N$ is parameter count, $D$ is the seen token count, and $P$ denotes the quantization bits. The factor $\mathrm{eff}(P)$ captures the effective capacity penalty due to quantization. $\mathrm{eff}(\text{FP})\equiv 1$.
We jointly fit $A,\alpha,B,\beta,E$ shared across methods and learn a separate $\mathrm{eff}(P)$ per method/bit-width via nonlinear least squares on the grid of $(N,D)$ we trained (Appendix B). We regularize with weak log-priors to keep $\alpha,\beta>0$. Figure 2 plots $\mathrm{eff}(P)$. CAGE consistently increases $\mathrm{eff}(P)$ vs. QuEST, with the largest gains at $b\approx 3,4$, indicating improved effective capacity for a given parameter budget.
Figure 3: MXFP4 QAT on Llama-3.2-3B (Tulu-SFT): CAGE vs. MXFP4 baseline, with/without Hadamard. CAGE improves GSM8K/HellaSwag/WinoGrande, with larger gains when Hadamard is enabled.
QAT fine-tuning on MXFP4.
We further evaluate CAGE under an MXFP4-style 4-bit floating format [9] on a larger model, Llama-3.2-3B [10]. We fine-tune on Tulu-SFT [19] with QAT (master FP32, forward W4A4 in MXFP4) with and without CAGE applied, and report zero-shot or few-shot scores on GSM8K (exact-match), HellaSwag (accuracy), and WinoGrande (accuracy). We test with and without the Hadamard pre-transform to isolate its interaction with MXFP4. Figure 3 summarizes results. Hyperparameters for the experiments can be found in Appendix B
Figure 4: Quadratic with non-isotropic Hessian: final validation loss across various condition numbers. CAGE reduces stationary error.
Figure 5: Representative trajectories (PC1/PC2) for the quadratic task; CAGE follows curvature and converges closest to the optimum.
Quadratic experiments.
[…]
5 Conclusion and Discussion

### v2

@@ 3.2 Theoretical Analysis
$x_{t+1}=x_{t}-\alpha(\widetilde{\nabla}f(x_{t})+\lambda(x_{t}-Q(x_{t})))$ (3)
for $t=0,1,\dots$. We make the following assumptions regarding the smoothness of the loss and the nature of the stochastic noise.
@@
Both assumptions are quite standard in the optimization literature and, in some sense, minimal to get meaningful convergence guarantees. Additionally, to handle the quantization operator $Q$ in the analysis, it is common to approximate the actual non-smooth quantization operator $Q$ with smooth surrogates with annealing hyperparameters.
@@
$\mathbb{E}\|\nabla_{\rm\lambda P}f(Q(\hat{x}))\|^{2}\leq\frac{2}{\sqrt{T}}\left(A_{1}+A_{2}+A_{3}\right)\max\left(1,\tfrac{L}{\sqrt{T}}\right),$
where $\hat{x}$ is a random iterate drawn from the history $\{x_{0},x_{1},\dots,x_{T-1}\}$ uniformly at random and
$A_{1}=f(x_{0})-f^{*},\;A_{2}=\frac{L\sigma^{2}}{2}$
$A_{3}=\lambda\max_{x\in[x_{0},x_{T}]}\|x-Q(x)\|\|x_{0}-x_{T}\|.$
@@ Discussion.
First, observe that the convergence bound shows strong ergodic convergence to a Pareto-optimal state of the problem in Equation 1, with a convergence rate that is optimal in the unquantized case. The convergence measure is defined with respect to our proposed Pareto-gradients (Equation (2)) and, importantly, there is no non-vanishing term in the upper bound because of quantization error. In practical settings, both the distance $\|x_{0}-x_{T}\|$ and the quantization error $\|x-Q(x)\|$ are expected to remain bounded for any total number of training steps $T$. Therefore, this implies that the iterates in Equation (3) achieve $\mathcal{O}(\frac{1}{\sqrt{T}})$ ergodic convergence to the Pareto-optimal solution.
@@ 3.3 Practical Implementation
Since our main application is in training and fine-tuning LLMs, we implement CAGE on top of AdamW loshchilov2018adamw. In contrast to LOTION kwun2025lotion, we focus on the decoupled update rule and keep the correction term outside the preconditioning path (Alg. 1). We use this choice in most experiments, as we have found it to work better in practice. The base optimizer AdamW processes the mini-batch gradients $g_{t}$ obtained from STE, and the CAGE correction acts as a lightweight, elementwise post-step on the parameters. This decoupling has multiple practical advantages: (1) it preserves the optimizer’s well-understood behavior, (2) it keeps the Pareto stationarity direction xQ(x)x-Q(x) from being distorted by preconditioning, which empirically stabilizes training in low-bit regimes, and (3) it avoids numerical errors by decoupling terms with different behavior.
@@
4:Initialize $m_{0}\leftarrow 0,\;v_{0}\leftarrow 0,\;e_{0}\leftarrow 0$
@@
10:  if $r_{t}>s$ then
11:   $\lambda_{t}\leftarrow\lambda\cdot\dfrac{r_{t}-s}{1-s}$
12:  end if
13:  Sample minibatch and compute stochastic gradient $g_{t}$ with quantized forward pass.
14:  $x_{t}\leftarrow(1-\alpha\omega)\,x_{t}$ $\triangleright$ decoupled weight decay
15:  $m_{t}\leftarrow\beta_{1}m_{t-1}+(1-\beta_{1})g_{t}$
16:  $v_{t}\leftarrow\beta_{2}v_{t-1}+(1-\beta_{2})\,g_{t}\odot g_{t}$
17:  $\hat{m}_{t}\leftarrow m_{t}/(1-\beta_{1}^{t})$; $\hat{v}_{t}\leftarrow v_{t}/(1-\beta_{2}^{t})$
18:  $\tilde{x}_{t+1}\leftarrow x_{t}-\alpha\,\hat{m}_{t}/\left(\sqrt{\hat{v}_{t}}+\varepsilon\right)$
19:  $e_{t}\leftarrow x_{t}-Q(x_{t})$ $\triangleright$ quantization error (no grad)
20:  $x_{t+1}\leftarrow\tilde{x}_{t+1}-\alpha\,\lambda_{t}\,e_{t}$ $\triangleright$ decoupled correction
21:end for
22:return $x_{T}$
@@
$\tilde{x}_{t+1}=x_{t}-\alpha\cdot\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\varepsilon},\text{ with }\hat{m}_{t}=\frac{m_{t}}{1-\beta_{1}^{t}},\;\hat{v}_{t}=\frac{v_{t}}{1-\beta_{2}^{t}}$
@@
$e_{t}\coloneq x_{t}-Q(x_{t}),\qquad x_{t+1}=\tilde{x}_{t+1}-\alpha\lambda_{t}\bar{e}_{t}.$
@@
$r_{t}\coloneq\frac{t}{T},\qquad\lambda_{t}=\begin{cases}0,&r_{t}\leq s,\\ \lambda\cdot\dfrac{r_{t}-s}{1-s},&r_{t}>s.\end{cases}$
Figure 1: Quadratic with non-isotropic Hessian: final validation loss for all methods various condition numbers (lower is better). The loss for CAGE does not include the regularizer term. We find that CAGE consistently reduces stationary error.
Figure 2: Representative trajectories (projected to the top two principal components PC1 and PC2) for the quadratic task; CAGE follows curvature and converges closest to the optimum.
In practice, we have found the parameters $s\in[0.8,0.95]$ with a linear ramp and $\lambda\in[1,2.5]$ to work robustly for AdamW. We provide sweeps justifying these choices in Appendix C. The ramp prevents abrupt shifts in the optimization objective and reduces loss spikes in training.
@@ 4 Experiments
Unless otherwise stated, our experiments use a standard row-wise matrix quantizer, e.g. qlora. Specifically, our QAT instantiation is based on QuEST panferov2025quest. This works as follows: let $x\in\mathbb{R}^{d}$ denote a tensor row, and let $H\in\{\pm\tfrac{1}{\sqrt{d}}\}^{d\times d}$ be a Hadamard transform of corresponding size. Before quantization, each tensor is “rotated” via the orthogonal Hadamard matrix:
@@
Given the target bit-width $b$, we quantize symmetrically with integer bounds $q_{\max}=2^{b-1}-1$ and $q_{\min}=-2^{b-1}$. Using the pre-computed MSE-optimal Gaussian clipping factor $k_{b}$ (per bit-width), with $\sigma=\sqrt{\tfrac{1}{d}\sum_{i}z_{i}^{2}}$, we set the scale
$s=\frac{k_{b}\,\sigma}{q_{\max}}.$
We then apply the quantization operation, rescale, and invert the transform:
@@
Note that CAGE itself is quantizer-agnostic. It augments the optimizer update via the instantaneous quantization error $e_{t}=x_{t}-Q(x_{t})$ (see §3.3). We choose the QuEST quantizer here because it is a strong SOTA baseline for low-bit training. We note that the QuEST paper already showed superior results in terms of both training stability and loss, relative to prior methods such as LSQ/LSQ+ esser2020lsq; bhalgat2020lsqplus. As such, we omit comparisons with further prior work, while noting that other quantizers can be used with CAGE without changing the method.
4.1 Synthetic Loss Experiments
To examine the exact QAT dynamics, we first study different methods optimizing a quadratic objective $f(x)=\tfrac{1}{2}x^{\top}Ax-b^{\top}x$ with $\kappa(A)\in\{1,10,100\}$ (non-isotropic curvature) over a quantized domain. We compare STE-SGD, STE-Adam, and CAGE-Adam (decoupled) under a 4-bit quantization of $x$, using the same pipeline as above, under different condition numbers.
[…]
Figure 3: QAT accuracy on Llama-3.2-3B (Tulu-SFT) for CAGE vs. the state-of-the-art MXFP4 baseline, using QuEST panferov2025quest. CAGE consistently improves GSM8K/HellaSwag/WinoGrande results, essentially halving the error due to quantization. Hadamard transform is applied in both cases for outlier mitigation.
Figure 4: Validation loss during QAT (W4A4) for AdamW, Shampoo, SOAP, and Muon, for standard QAT and CAGE. Solid lines denote training with CAGE, dashed – without. Across all optimizers, the method yields lower validation loss.
4.2 Quantization-Aware Fine-tuning
Next, we evaluate the suitability of CAGE in a more realistic setting, where it is used to perform QAT over the MXFP4 4-bit floating format mxspec on the Llama-3.2-3B model llama3. We fine-tune on Tulu-SFT lambert2024tulu3 with QAT (master weights in FP32, the forward pass has both weights and activations quantized to MXFP4), following the recently-proposed QAT recipe of egiazarian2025bridginggappromiseperformance, which yields state-of-the-art results for MXFP4.
We report zero-shot or few-shot scores on GSM8K (exact-match), HellaSwag (accuracy), and WinoGrande (accuracy). (Higher is better.) Hyperparameters for the experiments are in Appendix B, where we also include the results without the Hadamard transform, which show similar trends.
Figure 4 summarizes the results. We observe that CAGE consistently improves accuracy relative to the state-of-the-art QAT baseline; specifically, it roughly halves the quantization error when mapping the model to MXFP4 in this format.
Figure 5: Validation loss versus model size (bytes) for weight and activation quantization using W2A2, W3A3 and W4A4, comparing CAGE+QuEST to baseline QuEST and BF16. Observe that CAGE yields consistently lower loss.
Figure 6: Fitted parameter efficiency $\mathrm{eff}(P)$, normalized relative to standard 16-bit, across precisions. CAGE improves effective model capacity over QuEST, peaking near 4-bit.
@@ 4.2 Quantization-Aware Fine-tuning
Method 30M 50M 100M 200M 430M 800M $\mathrm{eff}(P)$
CAGE + HT 26.277 22.747 18.944 16.166 13.789 12.182 0.797
QuEST + HT 26.475 23.062 19.123 16.311 14.169 12.482 0.733
CAGE (no HT) 27.287 23.781 19.630 16.596 14.024 12.341 0.705
QuEST (no HT) 27.401 23.991 19.799 17.093 14.375 12.797 0.620
BF16 (reference) 24.715 21.491 17.923 15.422 13.176 11.698 1.0
4.3 Pretraining Experiments
Next, we test our method by pre-training Llama-style transformers with parameter counts $N\in\{30\text{M},50\text{M},100\text{M},200\text{M},430\text{M},800\text{M}\}$ under quantization-aware training from scratch, with weight/activation bit-widths $b\in\{2,3,4\}$. Our baseline is again the QuEST QAT approach, which uses a Hadamard transform prior to quantization panferov2025quest, which was shown to be superior to classic methods such as STE and LSQ esser2020lsq. Training uses the C4 dataset (c4) with a fixed budget of $D=100\times N$ tokens, for a token-per-parameter (TPP) ratio of 100, where $D$ is the number of tokens and $N$ is the number of model parameters. Master weights are stored FP32, and updates follow AdamW with decoupled weight decay loshchilov2018adamw as in Algorithm 1.
We compare CAGE against (i) QuEST (the current SOTA baseline) and (ii) a high-precision BF16 training reference. Figure 6 shows validation loss vs. model size for $b\in\{2,3,4\}$, with CAGE, QuEST baseline, and BF16.
Experiments are done on 8xH100 GPUs using standard hyperparameters (LR schedule, warmup, clipping, etc.), presented in Appendix B. Each pretraining result is an average over three seeds and the mean and standard deviation are reported in Table 1. Models up to $200\text{M}$ are visualized in Figure 6. We note that the standard deviation is too small to be visible in Figure 6.
Loss comparison.
Table 1 presents the actual validation loss results for W4A4 training across methods and model sizes. We observe that CAGE consistently improves upon QuEST, by significant margins, across all model sizes, and both with and without the Hadamard transform.
Figure 6 examines the validation loss-vs-model-size Pareto front induced by CAGE and QuEST across W4A4, W3A3, and W2A2 precisions. We observe that CAGE is consistently Pareto-superior relative to QuEST, at every precision. W4A4 CAGE appears to be Pareto-optimal in terms of obtaining the lowest loss at a fixed model size, with a small advantage relative to W3A3. Notably, W3A3 CAGE is Pareto-superior to W4A4 QuEST!
Loss vs. Regularizer.
In Figure 7, we visualize the evolution of the loss in parallel to the quantization error for pre-training the 50M Llama model in W4A4. We observe that, in the case of the baseline, there is a clear increase in quantization error as we continue training; at the same time, there is a sharp, consistent decrease in quantization error once the regularization starts, proportional to the value of the $\lambda$ parameter. This confirms the predicted behavior of CAGE, as well as the improvements left “on the table” by the standard QAT baseline.
Universality Across Optimizers.
To verify that gains produced by CAGE are independent of the optimizer update, we conduct a series of additional experiments using alternative optimizers to AdamW: Muon muon, Shampoo shampoo, and SOAP soap.
We train 50M models from the OLMo2 family olmo2 on the ClimbMix dataset climb_dataset for $\frac{D}{N}=100$ training tokens per parameter. OLMo2 models use no biases, employ rotary positional embeddings rope, RMSNorm rmsnorm, and reordered pre-normalization swin; olmo2. For each optimizer, we first tune the hyperparameters, and then we match the training setup and weights initialization to compare training curves with and without employing CAGE method. We use CAGE hyperparameters $s=0.9$, $\lambda=5$.
As shown in Figure 4, CAGE consistently reduces validation loss for all examined optimizers without any additional hyperparameter tuning. Since improvements introduced by CAGE do not depend on the choice of optimizer, we conclude that it can be regarded as an optimizer-agnostic technique.
Figure 7: Illustration of the validation loss for training the 50M Llama model in W4A4 using the QuEST baseline end-to-end (left), focused on the last fraction of steps (middle), and in terms of quantization error (regularizer) values (right). Observe that, while the overall training behavior is similar, CAGE significantly reduces loss for the quantized model in the final part of training.
Precision Scaling Law.
To compare methods beyond pointwise losses, we fit a separable data/model/precision-scaling law kumar2024scaling; frantar2025compressionscaling to the validation loss. Following references kumar2024scaling; panferov2025quest; hoffmann2022trainingcomputeoptimallargelanguage; frantar2025compressionscaling, we fit a law of the form:
@@ 4.3 Pretraining Experiments
where $N$ is parameter count, $D$ is the seen token count, and $P$ denotes the bit budget. The factor $\mathrm{eff}(P)$ captures the effective capacity penalty due to quantization, where the capacity of standard precision is $\mathrm{eff}(\text{FP})=1$.
To fit the law, we jointly fit $A,\alpha,B,\beta,E$ shared (jointly) across methods and learn a separate $\mathrm{eff}(P)$ per method/bit-width via nonlinear least squares on the grid of $(N,D)$ we trained (see details in Appendix B). We regularize with weak log-priors to maintain shared values for the parameters $\alpha,\beta>0$ across precisions.
Improved Parameter Efficiency.
The obtained $\mathrm{eff}(P)$, normalized relative to standard 16-bit precision by multiplying by $16/P$, are plotted in Figure 6. We observe that CAGE consistently increases $\mathrm{eff}(P)$ vs. QuEST (and the other baselines), indicating improved effective capacity for a given parameter budget. Specifically, we find that CAGE improves parameter efficiency by more than 10% at 4-bit precision, and 20% at 2-bit precision, relative to the prior state-of-the-art method. We also observe that 4-bit precision appears to be the “optimal” bit-width, although only with a small relative advantage relative to 3-bit when using CAGE.
Comparison to LOTION.
We also compare CAGE to LOTION. At the time of writing, no official implementation of LOTION was publicly available, so we reimplemented the method based on the description in the paper. We performed an extensive sweep over the regularization coefficient and found that the best-performing setting for LOTION corresponds to a relatively large value (in the order of $4\times 10^{3}$). Since LOTION assumes weight-only quantization and is sensitive to saturation, we used an AbsMax weight quantizer (W4A16) in this experiment, without activation quantization.
The comparison was carried out in our standard pretraining setup: a 100M-parameter Llama-style model trained on C4 with QAT. As shown in Figure 8, both methods track a similar loss curve in the early stages of training, but CAGE achieves a clearly lower final validation loss, while LOTION plateaus. We note that we will re-run this comparison once an official LOTION implementation becomes available, to rule out discrepancies due to reimplementation.
Figure 8: Pretraining loss comparison between CAGE and LOTION on a 100M-parameter Llama-style model trained on C4 under weight-only QAT with AbsMax W4A16 quantization. Both methods follow a similar trajectory early in training, but CAGE continues to improve and attains a substantially lower final loss, while LOTION plateaus.
5 Summary and Discussion